# Safeguarding Blockchain Ecosystem: Understanding and Detecting Attack Transactions on Cross-chain Bridges

### v1

@@ Abstract.
Cross-chain bridges are essential decentralized applications (DApps) to facilitate interoperability between different blockchain networks. Unlike regular DApps, the functionality of cross-chain bridges relies on the collaboration of information both on and off the chain, which exposes them to a wider risk of attacks. According to our statistics, attacks on cross-chain bridges have resulted in losses of nearly 4.3 billion dollars since 2021. Therefore, it is particularly necessary to understand and detect attacks on cross-chain bridges. In this paper, we collect the largest number of cross-chain bridge attack incidents to date, including 49 attacks that occurred between June 2021 and September 2024. Our analysis reveal that attacks against cross-chain business logic cause significantly more damage than those that do not. These cross-chain attacks exhibit different patterns compared to normal transactions in terms of call structure, which effectively indicates potential attack behaviors. Given the significant losses in these cases and the scarcity of related research, this paper aims to detect attacks against cross-chain business logic, and propose the BridgeGuard tool. Specifically, BridgeGuard models cross-chain transactions from a graph perspective, and employs a two-stage detection framework comprising global and local graph mining to identify attack patterns in cross-chain transactions. We conduct multiple experiments on the datasets with 203 attack transactions and 40,000 normal cross-chain transactions. The results show that BridgeGuard’s reported recall score is 36.32% higher than that of state-of-the-art tools and can detect unknown attack transactions.
@@
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXccs: Security and privacy Web application securityccs: Applied computing Electronic funds transfer
@@ 1. Introduction
Cross-chain bridges, through smart contracts and other technical means, enables users to swiftly transfer assets between different blockchain networks, thus achieving cross-chain liquidity of assets. According to DappRadar [^1] https://dappradar.com/rankings/defi/18?category=defi_cross-chain, there are currently over 440 cross-chain bridge DApps implemented based on various cross-chain mechanisms, making them an indispensable part of the blockchain ecosystem. For example, Celer cBridge operates based on the Hash Time Lock Contract (HTLC) mechanism (Yu et al., 2019), Poly Network operates based on Relay Chain (Singh et al., 2020), and MultiChain operates based on Notary Mechanism (Xiong et al., 2022). With the increasing number of blockchain networks, the importance and demand for cross-chain bridges are gradually becoming more prominent.
However, as bridges between different blockchain networks, cross-chain bridges often carry a substantial amount of asset value, and their hidden vulnerabilities make them targets for hackers. In recent years, many security incidents of cross-chain bridges have emerged. The top three security incidents with the highest losses in the Rekt attack database [^2] https://rekt.news/zh/leaderboard/ are all related to cross-chain bridges, with Ronin Network losing \$624 million, Poly Network losing \$611 million, and BNB Bridge losing \$586 million, respectively. Particularly, Thorchain suffered three attacks within just two months (on June 29, July 16, and July 23, 2021). The frequency of these security incidents indicates that the security issues of cross-chain bridges have become a significant challenge in the current blockchain field.
Despite the attention paid to the safety of cross-chain bridges, relatively few studies and solutions have been developed in this area. Although some studies have explored and analyzed the safety issues of cross-chain bridges, there are still some limitations. For example, a recent work Xscope (Zhang et al., 2022) focuses on investigating and summarising the security incidents of cross-link bridges occurring from 2021 to March 2022 and gives a rule-based detection method. And some Systematisation of Knowledge Papers (SoK) studies  (Notland et al., 2024a; Lee et al., 2023a, b) classify and discuss the attack surface, defense methods and problems of cross-chain bridges. However, as cross-chain mechanisms are still under development and refinement, and there are various attacks against cross-chain bridges. There is still a lack of comprehensive analyses of cross-chain bridge attacks as well as scalable solutions.
@@ 2.2.1. Data Collection and Statistic
Finally, we collect 49 cross-chain bridge attack incidents that occurred between June 2021 and September 2024, which is the largest academic dataset as far as we know. The comprehensive list of cross-chain bridge attack incidents is shown in the Appendix A.3, which includes details such as the attacked cross-chain bridges, attack date, the amount of losses, information source, attack stage of cross-chain, and reasons.
Based on the cross-chain business logic introduced in Section 2.1, we first categorize the collected attack incidents as either against or not against cross-chain business logic. First, we investigate the incidents not against cross-chain business logic. These incidents include private key leaks, flash loans, rug pulls, front-end hacking, etc. See more details in Appendix A.3. Then, we analyze the number of incidents and the corresponding financial losses. As shown in Fig. 2, out of the 49 incidents we collected, 27 did not against cross-chain business logic. However, the financial losses caused by attacks against cross-chain business logic were nearly six times greater than those from non-cross-chain business logic attacks.
@@ 2.3.1. Attack on Source Chain (denoted as $\mathcal{A}_{src}$)
Attacks on Router Contracts. Attacks against router contracts are built on the basis that there is an existing token lock, but the business logic of the router contract is faulty, thus generating a fake deposit event. Take ThorChain #1 as an example. In the ThorChain #1 incident [^7] https://hacked.slowmist.io/zh/?c=Bridge, the attacker performed a token lock, but the token’s ERC20 token symbol is “ETH”. However, there is a logic error in the router contract that recognizes the tokens that are topped up as genuine Ether ETH.
@@ 2.3.2. Attack Off Chain
Most off-chain attacks against cross-chain Bridges are aimed at external authentication. This may be because the bridges will choose an external verification mechanism to achieve fast multi-chain adaptation. There exists an impossible triangle for cross-chain interoperability, which means that any cross-chain scheme design cannot balance scalability, no need for trust, and easy adaptation (Arjun Bhuptani, 2021). While external validation enables fast multi-chain adaptation, it introduces new trust assumptions. Therefore, the external verification approach is one of the more fragile of all cross-chain mechanisms.
@@ 2.4. Discussion of Attack Transactions
3. BridgeGuard: Detecting Cross-chain Attacks Transactions
@@ 3.2.2. Global Graph Mining of xTEG
Graph embedding techniques can effectively transform high-dimensional discrete graph data into low-dimensional continuous vector spaces, maximally preserving the structural properties of the graph (Xu, 2021). Therefore, We employ a graph embedding method Graph2vec (Narayanan et al., 2017) to learn global features in xTEG with the following key advantages: unsupervised representation learning that captures structural equivalence, i.e., structurally similar graphs can produce similar embeddings. Graph2vec algorithm extends the concept of word embedding from the Doc2vec algorithm (Lau and Baldwin, 2016) in natural language processing to graph embedding. It treats the entire graph as a document and considers each vertex’s rooted subgraph (i.e., neighborhood) as the words in the document. The basic idea of using Graph2vec for xTEG graph mining is as follows: for each vertex in the xTEG, it first generates rooted subgraphs using the Weisfeiler-Lehman kernel (WL kernel) (Shervashidze et al., 2011) and assigns unique labels to these subgraphs. Then, it treats the collection of all rooted subgraphs around each vertex as its vocabulary. Finally, it employs the Skip-gram optimization model (Guthrie et al., 2006) from Doc2vec to learn vector representations for each xTEG in the dataset.
@@ 3.3. Experimental Setup
In this part, we perform experiments to demonstrate the effectiveness of the proposed tool, BridgeGuard, in protesting cross-chain bridges against attacks.
Research questions. In particular, we aim to answer the following research questions (RQs):
@@
Dataset. Based on the data collection method introduced in Section 2.2, we collect 49 incidents occurred on cross-chain bridges from June 2021 and September 2024. Given that the number of normal transactions on the chain far exceeds the number of attack transactions, in order to make the distribution of the evaluation dataset as close as possible to the actual situation, we mix 203 attack transactions into 40,000 normal transactions at a rate of 5%. In the supervised tasks, we divide the dataset into a training set and a test set with a ratio of 7:3. Besides, we conduct ten repeated experiments to obtain averaged results.
@@ 3.5. RQ2: Comparison with Existing tools
Table 4. Results of different tools in detecting cross-chain attack transactions
@@ 3.6. RQ3: Finding New Attack Transaction
  • Attack transactions in Thorchain #1 incident: Two new attack transactions, 0x99f and 0x152 are detected. The sender of these transactions is the same as that of the reported attack transaction1010100x92b466c1908571c45b1a4e751550877a46c5f2ecbc308e01242ec6d2013ad88c. Based on the findings of Su et al. (Su et al., 2021), transactions initiated by the attacker are highly likely to be attack transactions as well. We also examine the behavior of these transactions, and find that the traces and triggered functions exhibited similar patterns to the known attack transaction.
@@
Table 5. Newly detected attack transactions by BridgeGuard
@@ 4.1. Security Analysis of Cross-chain Bridges
Lee et al. (Lee et al., 2023b) elucidated several cross-chain bridging attacks and proposed mitigations for most of them. Notlandet al. (Notland et al., 2024b) analyzed 34 cross-chain bridge security incidents, identifying 8 categories of critical vulnerabilities and proposing 11 mitigations. However, these studies are still insufficient in the systematic and comprehensive analysis of attacks, and may not cover all potential attack vectors. Belchioret al. (Belchior et al., 2023) proposed the Hephaestus model which provides a new way of modeling the complexity of cross-chain applications, but its applicability in real-world environments has yet to be verified. Zhanget al. (Zhang et al., 2022) discovered three types of vulnerabilities in cross-chain bridges and proposed the Xscope monitoring tool, but its validity and extensibility still require However, its effectiveness and scalability still need to be further studied. Therefore, future research should focus on integrating the existing results and conducting more systematic empirical analysis to improve the security and reliability of cross-chain bridges.
@@ 4.2. Detection for DeFi Attacks
Research on DeFi attacks can be divided into two types: detecting from a contract perspective and detecting from a transaction perspective. From the perspective of contracts, Rodler et al. (Rodler et al., 2018) mainly used the execution flow analysis method to detect re-entry vulnerabilities in contracts. And Chen  et al. (Chen et al., 2020) developed a tool that can detect contract security online and expand to custom vulnerabilities. From the perspective of transactions, Zhou et al. (Zhou et al., 2020) conducted a large-scale measurement and analysis of Ethereum transaction logs for the first time and discovered some new types of attacks, such as airdrop hunting. However, Su et al. (Su et al., 2021) focused on existing attack cases and proposed the tool DEFIER to automatically investigate large-scale attack events. Zhang et al. (Zhang et al., 2020) hoped to develop a universal attack detection framework, which detects the security of transactions by modifying Geth, replaying historical transactions, and defining a series of security attributes. In addition, Zhou et al. (Zhou et al., 2023) studied how to systematically measure, evaluate, and compare DeFi attack events. The paper (Wang et al., 2022) focuses on the detection of logical vulnerabilities on Ethereum. Su  et al. (Su et al., 2023) analyzed token leakage vulnerabilities by mining the relationship between users and DApps.
@@ 5. Conclusion and Future Work
In this paper, we conducted an in-depth study of cross-chain bridge attack incidents and proposed a detection tool for attacks targeting cross-chain business processes, called BridgeGuard. By collecting and analyzing 49 cross-chain bridge attack incidents, particularly those against cross-chain business processes, we constructed cross-chain transaction execution graphs (xTEGs) and extracted statistical and structural features. Experimental results show that BridgeGuard demonstrates excellent performance in detecting cross-chain attacks, with a recall rate 42.5% higher than the state-of-the-art tools and the ability to identify newly discovered attack transactions. We believe that the introduction of BridgeGuard provides an effective solution to enhance the security of cross-chain bridges, while also serving as an important reference for future research in cross-chain bridge security.
For future work, we plan to explore several directions. Firstly, we wish extend BridgeGuard to other types of cross-chain bridges, such as NFT bridges and governance bridges, to achieve more comprehensive cross-chain security monitoring. Additionally, we can optimize the performance of BridgeGuard, including improving detection efficiency and reducing resource consumption, to meet the requirements of real-world applications. Finally, we can explore the application of large-scale language model (LLM) in cross-chain security to improve the recognition and defense against complex attack patterns.
@@ Appendix A Appendix
Harmony 2022/06/23 100000000 Rekt News Not specific to cross-chain process Private key leakage
@@
Zenon 2021/11/21 1000000 Rekt Not specific to cross-chain process Flash-loan
@@
A.1. Background
A.1.1. Blockchain
Blockchain technology was first introduced by Satoshi Nakamoto in the Bitcoin whitepaper in 2008, representing a distributed and tamper-resistant ledger system designed to address trust among multiple parties in a public ledger  (Zheng et al., 2018). This technology operates on a peer-to-peer network, where each participant or node holds a copy of the entire blockchain. By leveraging cryptographic techniques, blockchain packages transactions into blocks and links them together in a chain, ensuring the security and transparency of data. Each block contains the hash value of the previous block, providing the blockchain with immutability and enabling trusted transactions without the need for intermediaries. Blockchains can be categorized into public chains, private chains, consortium chains, and others.
The emergence of multi-chain ecosystems has made it challenging for assets and data to interoperate between different blockchain networks. Currently, the blockchain ecosystem consists of multiple chains, with records of 260 public blockchains as of March 2024, according to data from DeFiLlama [^11] https://defillama.com/chains. However, data between different blockchain systems are not interoperable, akin to isolated islands. Therefore, DeFi bridges, as applications capable of facilitating asset circulation and information exchange between chains, can promote further development of the multi-chain ecosystem (Lee et al., 2023a).
A.1.2. Transaction and Smart Contract
Transaction data is a type of data on the blockchain. Transactions are data structures that record cryptocurrency information on the blockchain, which can be messages sent to smart contracts or simple token transfers to blockchain users. Transactions are fundamental units of activity on the blockchain, representing a modification to the blockchain’s state. In blockchains like Bitcoin and Ethereum, transactions typically include sender addresses, recipient addresses, amounts of assets transferred, transaction fees, and other information. Once a transaction is created, it is broadcasted to all nodes on the network, undergoes validation, and is included in a new block. Transactions on the blockchain are irreversible; once confirmed, they are permanently recorded on the blockchain.
Smart contracts are another data type in the blockchain domain, initially invented in Ethereum. Smart contracts are self-executing contracts that run on the blockchain, where the terms and conditions are programmatically defined and executed by the blockchain network (Zheng et al., 2020). Smart contracts are typically written in programming languages like Solidity and deployed onto the blockchain for execution. Once deployed on the blockchain, smart contracts become immutable. They automatically execute once their predefined conditions are met, without the need for third-party intervention.
A.2. Cross-chain Bridge Attack Incidents List

### v2

@@ Abstract.
Cross-chain bridges are essential decentralized applications (DApps) to facilitate interoperability between different blockchain networks. Unlike regular DApps, the functionality of cross-chain bridges relies on the collaboration of information both on and off the chain, which exposes them to a wider risk of attacks. According to our statistics, attacks on cross-chain bridges have resulted in losses of nearly \$4.3 billion since 2021. Therefore, it is particularly necessary to understand and detect attacks on cross-chain bridges. In this paper, we collect the largest number of cross-chain bridge attack incidents to date, including 49 attacks that occurred between June 2021 and September 2024, of which 22 were attacks on cross-chain bridge business logic. Our analysis reveal that attacks against cross-chain business logic cause significantly more damage than those that do not. These cross-chain attacks exhibit different patterns compared to normal transactions in terms of call structure, which effectively indicates potential attack behaviors. Given the significant losses in these cases and the scarcity of related research, this paper aims to detect attacks against cross-chain business logic, and propose the BridgeGuard tool. Specifically, BridgeGuard models cross-chain transactions from a graph perspective, and employs a two-stage detection framework comprising global and local graph mining to identify attack patterns in cross-chain transactions. We conduct multiple experiments on the datasets with 203 attack transactions and 40,000 normal cross-chain transactions. The results show that BridgeGuard’s reported recall score is 36.32% higher than that of state-of-the-art tools and can detect unknown attack transactions.
@@
journalyear: 2025copyright: acmlicensedconference: Proceedings of the ACM Web Conference 2025; April 28-May 2, 2025; Sydney, NSW, Australiabooktitle: Proceedings of the ACM Web Conference 2025 (WWW ’25), April 28-May 2, 2025, Sydney, NSW, Australiadoi: 10.1145/3696410.3714604isbn: 979-8-4007-1274-6/25/04ccs: Security and privacy Web application securityccs: Applied computing Electronic funds transfer
@@ 1. Introduction
Cross-chain bridges, through smart contracts and other technical means, enables users to swiftly transfer assets between different blockchain networks, thus achieving cross-chain liquidity of assets. According to DappRadar [^1] https://dappradar.com/rankings/defi/18?category=defi_cross-chain, there are currently over 440 cross-chain bridge DApps implemented based on various cross-chain mechanisms, making them an indispensable part of the blockchain ecosystem. However, as bridges between different blockchain networks, cross-chain bridges often carry a substantial amount of asset value, and their hidden vulnerabilities make them targets for hackers. In recent years, many security incidents of cross-chain bridges have emerged. The top three security incidents with the highest losses in the Rekt attack database [^2] https://rekt.news/zh/leaderboard/ are all related to cross-chain bridges, with Ronin Network losing \$624 million, Poly Network losing \$611 million, and BNB Bridge losing \$586 million, respectively. Particularly, Thorchain suffered three attacks within just two months (on June 29, July 16, and July 23, 2021). The frequency of these security incidents indicates that the security issues of cross-chain bridges have become a significant challenge in the current blockchain field.
Despite growing attention to cross-chain bridge safety, few studies and solutions have been developed. While some research, such as Xscope (Zhang et al., 2022), analyzes security incidents and proposes rule-based detection, and SoK papers (Notland et al., 2024a; Lee et al., 2023a, b) discuss attack surfaces and defense methods, there is still a lack of comprehensive analysis and scalable solutions due to the ongoing development of cross-chain mechanisms and the variety of attacks.
@@ 2.2.1. Data Collection and Statistic
Finally, we collected 49 cross-chain bridge attack incidents from June 2021 to September 2024, the largest academic dataset to date. The detailed list, including attack details, is in Appendix A.1. Based on the cross-chain business logic, we categorize the incidents into those targeting business logic and those not. The non-business logic incidents include private key leaks, flash loans, and rug pulls, detailed in Appendix A.2. Of the 49 incidents, 22 targeted cross-chain business logic, causing financial losses nearly six times greater than non-business logic attacks, as shown in Fig. 2.
@@ 2.3.1. Attack on Source Chain (denoted as $\mathcal{A}_{src}$)
Attacks on Router Contracts. Attacks against router contracts are built on the basis that there is an existing token lock, but the business logic of the router contract is faulty, thus generating a fake deposit event. Take ThorChain #1 as an example. In the ThorChain #1 incident [^7] https://hacked.slowmist.io/zh/?c=Bridge, the attacker performed a token lock, but the token’s ERC20 token symbol is “ETH”. However, there is a logic error in the router contract that recognizes the tokens that are topped up as genuine Ether.
@@ 2.3.2. Attack Off Chain
Most off-chain attacks against cross-chain Bridges are aimed at external authentication. This may be because the bridges will choose an external verification mechanism to achieve fast multi-chain adaptation (Arjun Bhuptani, 2021). While external validation enables fast multi-chain adaptation, it introduces new trust assumptions. Therefore, the external verification approach is one of the more fragile of all cross-chain mechanisms.
@@ 2.4. Discussion of Attack Transactions
3. BridgeGuard: Detecting Cross-chain Attack Transactions
@@ 3.2.2. Global Graph Mining of xTEG
Graph embedding techniques can effectively transform high-dimensional discrete graph data into low-dimensional continuous vector spaces, maximally preserving the structural properties of the graph (Xu, 2021). Therefore, We employ a graph embedding method Graph2vec (Narayanan et al., 2017) to learn global features in xTEG with the following key advantages: unsupervised representation learning that captures structural equivalence, i.e., structurally similar graphs can produce similar embeddings. Graph2vec algorithm extends the concept of word embedding from the Doc2vec algorithm (Lau and Baldwin, 2016) in natural language processing to graph embedding. It treats the entire graph as a document and considers each vertex’s rooted subgraph (i.e., neighborhood) as the words in the document. The basic idea of using Graph2vec for xTEG graph mining is as follows: for each vertex in the xTEG, it first generates rooted subgraphs using the Weisfeiler-Lehman kernel (WL kernel) (Shervashidze et al., 2011) and assigns unique labels to these subgraphs. Then, it treats the collection of all rooted subgraphs around each vertex as its vocabulary. Finally, it employs the Skip-gram optimization model (Guthrie et al., 2006) from Doc2vec to learn 16 dimensional vector representations for each xTEG in the dataset.
@@ 3.3. Experimental Setup
In this part, we perform experiments to demonstrate the effectiveness of the proposed tool, BridgeGuard, in protesting cross-chain bridges against attacks. In particular, we aim to answer the following research questions (RQs):
@@
Dataset. Based on the data collection method introduced in Section 2.2, we collect 22 attacks against cross chain business logic from June 2021 and September 2024. Given that the number of normal transactions on the chain far exceeds the number of attack transactions, in order to make the distribution of the evaluation dataset as close as possible to the actual situation, we mix 203 attack transactions into 40,000 normal transactions at a rate of 0.5%. In the supervised tasks, we divide the dataset into a training set and a test set with a ratio of 7:3. Besides, we conduct ten repeated experiments to obtain averaged results.
@@ 3.5. RQ2: Comparison with Existing tools
Table 4. Comparison with different tools
@@ 3.6. RQ3: Finding New Attack Transaction
  • Attack transactions in Thorchain #1 incident: Two new attack transactions, 0x99f and 0x152 are detected. The sender of these transactions is the same as that of the reported attack transaction 0x92b Based on the findings of Su et al. (Su et al., 2021), transactions initiated by the attacker are highly likely to be attack transactions as well. We also examine the behavior of these transactions, and find that the traces and triggered functions exhibited similar patterns to the known attack transaction.
@@
Table 5. Newly detected attack transactions
@@ 4.1. Security Analysis of Cross-chain Bridges
Lee et al. (Lee et al., 2023b) describe several cross-chain bridging attacks and propose preventive measures for most of them. Notlandet al. (Notland et al., 2024b) analyzed 34 cross-chain bridge security incidents, identified 8 categories of critical vulnerabilities, and proposed 11 preventive measures. However, these studies are still insufficient in the systematic and comprehensive analysis of attacks, and may not cover all potential attack vectors. Zhang et al. (Zhang et al., 2022) discovered three types of vulnerabilities in cross-chain bridges and proposed the Xscope monitoring tool, but its effectiveness and scalability still need to be further studied. Therefore, future research should focus on integrating the existing results and conducting more systematic empirical analysis to improve the security and reliability of cross-chain bridges.
@@ 4.2. Detection for DeFi Attacks
Research on DeFi attacks can be divided into two types: detecting from a contract perspective and detecting from a transaction perspective. From the perspective of contracts, Rodler et al. (Rodler et al., 2018) mainly used the execution flow analysis method to detect re-entry vulnerabilities in contracts. And Chen  et al. (Chen et al., 2020) developed a tool that can detect contract security online and expand to custom vulnerabilities. From the perspective of transactions, Zhou et al. (Zhou et al., 2020) conducted a large-scale measurement and analysis of Ethereum transaction logs for the first time and discovered some new types of attacks, such as airdrop hunting. However, Su et al. (Su et al., 2021) focused on existing attack cases and proposed the tool DEFIER to automatically investigate large-scale attack events.
@@ 5. Conclusion and Future Work
In this paper, we studied cross-chain bridge attack incidents and proposed BridgeGuard, a detection tool for attacks targeting cross-chain business processes. By analyzing 49 incidents, including 22 against business processes, we constructed cross-chain transaction execution graphs (xTEGs) and extracted features. Experimental results show that BridgeGuard outperforms state-of-the-art tools, with a recall rate 42.5% higher and the ability to detect newly discovered attack transactions. We believe BridgeGuard provides an effective solution for cross-chain bridge security and a valuable reference for future research.
For future work, we plan to explore several directions. Firstly, we aim to extend BridgeGuard to other types of cross-chain bridges, such as NFT and governance bridges, in order to achieve more comprehensive security monitoring. Secondly, we will focus on optimizing BridgeGuard’s performance by improving detection efficiency and reducing resource consumption, thus meeting the requirements of real-world applications. Finally, we intend to investigate the application of large-scale language models (LLM) in cross-chain security to enhance the recognition and defense against complex attack patterns.
Acknowledgements.
The work described in this paper is supported by the National Key Research and Development Program of China (2023YFB2704700), the National Natural Science Foundation of China (62372485, 623B2102 and 62332004), the Natural Science Foundation of Guangdong Province (2023A1515011314), the Fundamental Research Funds for the Central Universities of China (24lgqb018).
@@ Appendix A Appendix
Harmony 2022/06/23 100,000,000 Rekt News Not specific to cross-chain process Private key leakage
@@
Zenon 2021/11/21 1,000,000 Rekt Not specific to cross-chain process Flash-loan
@@
Table 7. A summary of the 27 attacks that were not against Cross-chain Bridge business logic.
Reason Number How BridgeGuard Can Be Used
Over-Authorisation 7 Over-authorisation is a vulnerability in cross-chain bridge contracts. If such vulnerable contracts are already deployed, BridgeGuard can assist detection using a pre-execution module.
Private Key Leakage 7 Private key leakage is often due to social engineering attacks. BridgeGuard cannot prevent such leaks, as crypto attack detection studies generally do not address this issue.
Flash-loan 2 By modeling transactions on-chain, BridgeGuard can detect such attacks (similar to the comparative method, DeFiScanner). However, flash-loan attacks are not the primary focus of this study.
Rug-pull 3 Not explicitly discussed in the study.
DNS Hijacking 3 DNS hijacking involves attackers redirecting users to fake domains or IPs and falls under network-layer attacks. These are outside the scope of application-layer attack detection, including BridgeGuard’s focus.
Others 5 These issues relate to ecosystem security but focus on social engineering, consensus, or network-layer attacks, not cross-chain logic. ( Phishing email: 1, False transaction: 1, Fabricated block: 1, Replay attack: 1, Hacked Twitter: 1)
A.1. Cross-chain Bridge Attack Incidents List